# Haptic bilateral teleoperation system for free-hand dental procedures

### v2

@@ III-A Eye-hand coordination controller
In a HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance. In teleoperated free-hand dental procedures, using an eye-in-hand camera is the most effective solution [43]. This approach enables the operator to maintain a consistently centered view of the operative field, offering a significant advantage over eye-to-hand camera configurations, where the view may be off-center with respect to the area of interest or obstructed by physical obstacles. In this context, the human operator is naturally guided by the video stream transmitted by the camera to provide commands to the FR via the HD. However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle. This creates a mismatch between the movements expected by the operator and those actually executed by the robot in the camera view, compromising the accuracy of the task.
The da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, thereby increasing the bulk of the equipment and the overall costs, provides the motion correspondence between the operator’s movements and those of the end-effector, observed in the camera view, by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [44]. In our eye-hand coordination controller, we provide the motion correspondence at the command level, by creating a coupling between the HD’s stylus and the robot’s TCP, and appropriately converting the motion of the former into the motion of the FR. To enhance the naturalness of the teleoperation, a rotational coupling is also established between the HD’s stylus and the FR’s TCP. This includes defining the desired relative rotation ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$, and performing manually the alignment. This is done by the human operator during what we indicate as the engagement phase.
The latter is presented in Section III-A1, while the human tremors filtering and the eye-hand coordination mapping, implementing the motion correspondence, are discussed in Section III-A2 and Section III-A3, respectively.
@@ III-A1 Engagement phase
${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left({\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}\right)^{\top},$ (4)
@@
where ${\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}$ is the fixed rotation between $\mathcal{H}_{b}$ and $\mathcal{R}_{b}$. We note that the position of the HD’s stylus is not relevant for coupling, so it is converted to the current position of the robot’s TCP. The obtained frame $\mathcal{R}_{ed}$ is real-time displayed in a visualization tool, together with the actual robot end-effector one $\mathcal{R}_{e}$, as shown in Figure 4, so that the human operator, by acting on $\mathcal{H}_{e}$, can align the frame $\mathcal{R}_{ed}$ with the frame $\mathcal{R}_{e}$. Once the alignment is accomplished, i.e.  $\left[{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]^{\top}{\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf I}_{3}$, where ${\bf I}_{3}\in SO(3)$ is the identity matrix, the engagement phase ends, i.e. HD and FR are aligned, and the human operator can begin teleoperation.
@@ III-A2 Human tremors filtering
At each control cycle, the EHCC retrieves the actual pose ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ of the HD’s stylus, corresponding to ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$. Exploiting such a pose to generate a command for the robot would result in transferring the human hand motion tremors, further accentuated by the control frequency. Therefore, ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is filtered online by a moving average filter for poses. At the $k$-th control cycle, the filtered position $\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is obtained from the cumulative sum ${\bf s}(k)$ of samples within a window of fixed size $n$:
$\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\frac{{\bf s}(k)}{m}$ (6)
where $m\leq n$ is the number of samples in the moving average window at the $k$-th control cycle, and the cumulative sum ${\bf s}(k)$ is:
${\bf s}(k)=\begin{cases}{\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k),&\text{if}~{}k<n;\\ {\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n),&\text{otherwise.}\end{cases}$ (7)
The filtered quaternion $\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is the eigenvector associated with the maximum eigenvalue of the ${\bf M}(k)$ matrix, of proper size [45], defined as:
${\bf M}(k)=\frac{{\bf c}(k)}{m},$ (8)
where ${\bf c}(k)$ is the cumulative sum of samples within a window of fixed size $n$, updated at each control cycle:
${\bf c}(k)=\begin{cases}{\bf c}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top},&\text{if}~{}k<n;\\ {\bf c}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}+&\\ \quad-{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\left[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\right]^{\top},&\text{otherwise.}\end{cases}$ (9)
@@ III-A3 Eye-hand coordination
In a HBTS, coupling two robots with different kinematics requires the use of mapping and scaling techniques. The literature shows that this coupling can be achieved by converting the motion of the HD into that of the FR through simple kinematic transformations, and appropriately scaling it to either expand the workspace size of the HD or enhance control accuracy [46]. In the eye-in-hand configuration, converting quantities between the two spaces without considering the viewing angle would lead to a loss of motion correspondence between the movements expected by the operator and those actually executed by the robot in the camera view. Therefore, in our EHCC, we transfer the HD’s twist to the FR through an eye-hand coordination mapping and introduce dynamic scaling factors to improve control accuracy based on the task requirements.
@@
and $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$ the relative rotation:
$\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}=\begin{bmatrix}r_{11}&r_{12}&r_{13}\\ r_{21}&r_{22}&r_{23}\\ r_{31}&r_{32}&r_{33}\end{bmatrix},$ (11)
obtained from the quaternion $\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}$:
$\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}=\left[\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1)\right]^{\top}\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).$ (12)
@@
${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=\begin{bmatrix}\dfrac{\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)}{T_{s}}\\[8.5359pt] \dfrac{\vartheta{\bf r}}{T_{s}}\end{bmatrix},$ (13)
where $T_{s}$ is the control period and ${\bf r}$ and $\vartheta$ are the axis and the angle corresponding to $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$, respectively:
$\vartheta$ $=\cos^{-1}{\left(\dfrac{r_{11}+r_{22}+r_{33}-1}{2}\right)},$ (14)
${\bf r}$ $=\frac{1}{2\sin{\vartheta}}\begin{bmatrix}r_{32}-r_{23}\\ r_{13}-r_{31}\\ r_{21}-r_{12}\end{bmatrix}.$ (15)
@@
$\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=&\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]{\boldsymbol{\Lambda}}\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]\\ &\cdot\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}{\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).\end{split}$ (16)
where $\otimes$ denotes the Kronecker product, ${\boldsymbol{\Lambda}}$ is the scaling diagonal matrix, of proper size, and ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)$ is:
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}{\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k)).$ (17)
We note that ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$ is the fixed rotation defining the mapping between quantities in $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$, and $\bf z$ is the axis of $\mathcal{H}_{e}$ mapped to the ${\bf o}$ axis of $\mathcal{R}_{e}$, according to ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$. The ${\bf o}$ axis represents the optical axis of the camera in $\mathcal{R}_{e}$, which we assume to coincide with the ${\bf z}$ axis of $\mathcal{R}_{e}$ for negligible values of the camera tilt angle. Therefore, a rotation of $\varphi(k)$ about the ${\bf z}$ axis of $\mathcal{H}_{e}$ results in a viewing angle of $\varphi(k)$, as shown in Figure 5.
[…]
Figure 5: Visualization of the rotational coupling between $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$.
The rotation matrix ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))$ ensures the motion correspondence between the operator’s movements and those of the FR’s TCP in the camera view, by appropriately rotating the desired velocities according to the viewing angle. By setting ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))={\bf I}_{3}$ the mapping between twists becomes viewing angle independent:
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}},\forall k.$ (18)
The desired FR’s TCP homogeneous transformation ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ is obtained by numerically integrating the desired twist in (16). As for the desired rotation ${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is obtained as:
@@
where ${\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)}$ is the rotation matrix corresponding to the axis $\frac{{\bf r}}{\left\lVert{\bf r}\right\rVert}$ and angle $\vartheta$:
${\bf r}$ $=\left({\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k-1)\right)^{\top}{\boldsymbol{\omega}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)T_{s},$ (20)
$\vartheta$ $=\left\lVert{\bf r}\right\rVert.$ (21)
@@
${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}.$ (23)
The eye-hand coordination mapping and scaling can be turned into an algorithmic procedure with steps provided in Algorithm 1.
@@
    $\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k):=&\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\right]\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\right]^{\top}{\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\end{split}$
@@
    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\left[{\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\right]{\boldsymbol{\Lambda}}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)$

### v3

@@ III-A Eye-hand coordination controller
In a HBTS, visual feedback plays a crucial role in overcoming the limitations caused by distance. In teleoperated free-hand dental procedures, using an eye-in-hand camera is the most effective solution [16]. This approach enables the operator to maintain a consistently centered view of the operative field, offering a significant advantage over eye-to-hand camera configurations, where the view may be off-center with respect to the area of interest or obstructed by physical obstacles. In this context, the human operator relies on the video of the camera to command the FR via the HD. However, in the eye-in-hand configuration, a rotation of the tool (unavoidable, to preserve dexterity) causes a rotation of the viewing angle, whose lack of compensation results in a loss of both motion and force correspondence between the movements and perceived interaction expected by the operator and those actually executed by the robot in the camera view. This highlights the importance of an appropriate coupling.
The literature shows that this coupling can be achieved through simple kinematic transformations and scaling strategies [44]. Common mapping strategies include position control, workspace drift control [11], and rate control [51] which define how the operator’s motions are translated into robot commands, complemented by force scaling [44] to map the interaction forces into the range of forces that the HD can exert. However, these methods generally neglect viewing angle effects, relying on display-level compensation and thus compromising force correspondence. Indeed, the da Vinci Surgical SystemTM, which requires the insertion of multiple tools to implement an eye-to-hand strategy, thereby increasing bulk and cost of the equipment, provides the motion correspondence by projecting the endoscope image above the surgeon’s hands by means of mirrored overlay optics [18]. In a similar manner, in eye-in-hand cases, motion correspondence is typically achieved by aligning the visual feedback with the operator’s control frame through online compensation of camera–control frame misalignment in the displayed view [64].
To overcome this limitation, we introduce an eye–hand coordination mapping preserving both motion and force correspondence. Within the proposed EHCC, this mapping is implemented through a view-consistent transfer of the HD twist to the FR and incorporates dynamic scaling factors to improve control accuracy according to task requirements. To enhance the naturalness of the teleoperation, a rotational coupling is also established between the HD’s stylus and the FR’s TCP. This includes defining the desired relative rotation ${\bf R}_{\mathcal{R}_{e}}^{\mathcal{H}_{e}}$, and aligning manually. This is done by the human operator during what we indicate as the engagement phase.
The latter is presented in Section III-A1, while the human tremors filtering and the eye-hand coordination mapping, implementing the motion correspondence, are discussed in Section III-A2 and Section III-A3, respectively. The eye-hand coordination implementing the force correspondence is described in Section III-C.
@@ III-A1 Engagement phase
${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ $={\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}{\bf R}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k){\bf R}_{\mathcal{R}_{e}}^{\mathcal{H}_{e}},$ (2)
@@
where ${\bf R}_{\mathcal{H}_{b}}^{\mathcal{R}_{b}}$ is the fixed rotation between $\mathcal{H}_{b}$ and $\mathcal{R}_{b}$. We note that the position of the HD’s stylus is not relevant for coupling, so it is converted to the current position of the robot’s TCP. The obtained frame $\mathcal{R}_{ed}$ is real-time displayed in a visualization tool, together with the actual robot end-effector one $\mathcal{R}_{e}$, as shown in Figure 4, so that the human operator, by acting on $\mathcal{H}_{e}$, can align the frame $\mathcal{R}_{ed}$ with the frame $\mathcal{R}_{e}$. Once the alignment is accomplished, i.e., ${\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k){\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)={\bf I}_{3}$, the engagement phase ends, i.e., HD and FR are aligned, and the human operator can begin teleoperation.
@@ III-A2 Human tremors filtering
At each control cycle, the EHCC retrieves the actual pose ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ of the HD’s stylus, corresponding to ${\bf T}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$. Exploiting such a pose to generate a command for the robot would result in transferring the human hand tremors, further accentuated by the control frequency. Therefore, ${\bf x}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is filtered online by a moving average filter for poses. At the $k$-th control cycle, the filtered position is obtained as $\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)={\bf s}(k)/m$, where $m\leq n$ is the number of collected samples at the $k$-th control cycle in the moving average window of size $n$, and ${\bf s}(k)$ is their cumulative sum:
${\bf s}(k)={\bf s}(k-1)+{\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)-\begin{cases}0,&\text{if}~k<n;\\ {\bf p}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n),&\text{otherwise.}\end{cases}$ (4)
The filtered quaternion $\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$ is the eigenvector associated with the maximum eigenvalue of ${\bf M}(k)={\bf C}(k)/m$ [36], where ${\bf C}(k)$ is the cumulative sum of samples within the window, updated at each control cycle:
${\bf C}(k)=\begin{cases}{\bf C}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigl[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigr]^{\top},&\text{if}~k<n;\\ {\bf C}(k-1)+{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigl[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\bigr]^{\top}+&\\ \quad-{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\bigl[{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-n)\bigr]^{\top},&\text{otherwise.}\end{cases}$ (5)
@@ III-A3 Eye-hand coordination
and $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$ the relative rotation obtained from
$\overline{\bf q}_{\mathcal{H}_{e}(k)}^{\mathcal{H}_{e}(k-1)}=\bigl[\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k-1)\bigr]^{\top}\overline{\bf q}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k).$ (7)
@@
${\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)=T_{s}^{-1}\begin{bmatrix}\Delta\overline{{\bf p}}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)^{\top},&\vartheta{\bf r}^{\top}\end{bmatrix}^{\top}$, (8)
where $T_{s}\in\mathbb{R}_{+}$ is the control period, and ${\bf r}\in\mathbb{R}^{3\times 1}$ and $\vartheta\in\mathbb{R}$ are the axis-angle representation corresponding to $\overline{{\bf R}}^{\mathcal{H}_{e}(k-1)}_{\mathcal{H}_{e}(k)}$, respectively.
@@
$\begin{split}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=&\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\bigr)\biggl\{{\boldsymbol{\Lambda}}\Bigl[\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\bigr)\\ \cdot&\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{b}}^{\mathcal{H}_{e}}(k)\bigr){\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)\Bigr]\biggr\}.\end{split}$ (9)
where $\otimes$ denotes the Kronecker product, ${\boldsymbol{\Lambda}}\in\mathbb{R}^{6\times 6}$ is the scaling diagonal matrix with ${\boldsymbol{\Lambda}}_{4:6,4:6}={\bf I}_{3}$, and
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\bf z}^{\mathcal{R}_{e}}\bigl(\varphi(k)\bigr){\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}.$ (10)
We note that ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$ is the fixed rotation defining the mapping between quantities in $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$, and $\bf z$ is the axis of $\mathcal{R}_{e}$, which we assume to coincide with the optical axis of the camera $\bf o$} in $\mathcal{R}_{e}$ for negligible values of the camera tilt angle. Therefore, a rotation of $\varphi(k)$ about the $\bf z$} axis of $\mathcal{R}_{e}$ results in a viewing angle of $\varphi(k)$. The $\bf z$ axis of $\mathcal{R}_{e}$ is mapped to the $\bf z$} axis of $\mathcal{H}_{e}$, according to ${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}$. Hence, in the absence of compliance along rotational axes, the same behavior is achieved by rotating of $\varphi(k)$ about the $\bf z$} axis of $\mathcal{H}_{e}$, as shown in Figure 5, and (10) can be equivalently written as:
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}{\bf R}_{\bf z}^{\mathcal{H}_{e}}\bigl(\varphi(k)\bigr).$ (11)
Figure 5: Visualization of the rotational coupling between $\mathcal{H}_{e}$ and $\mathcal{R}_{e}$: (top) top view; (bottom) user view; (left) $\varphi=0$; (right) $\varphi=\pi/2$.
The rotation ${\bf R}_{\bf z}^{\mathcal{R}_{e}}\bigl(\varphi(k)\bigr)$, or equivalently ${\bf R}_{\bf z}^{\mathcal{H}_{e}}\bigl(\varphi(k)\bigr)$, ensures the motion correspondence between the operator’s movements and those of the FR’s TCP in the camera view, by appropriately rotating the desired velocities according to the viewing angle. By setting ${\bf R}_{\bf z}^{\mathcal{R}_{e}}(\varphi(k))={\bf I}_{3}$ (or equivalently ${\bf R}_{\bf z}^{\mathcal{H}_{e}}(\varphi(k))={\bf I}_{3}$) the mapping between twists becomes viewing angle independent:
${\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)={\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}},\;\forall k\in\mathbb{N}_{0}.$ (12)
We note that the angle $\varphi(k)$ is obtained from the exact measurements provided by the FR’s forward kinematics (or equivalently by the HD’s). For this reason, even during prolonged procedures, no orientation drift is accumulated, and motion correspondence is consistently preserved.
The desired FR’s TCP homogeneous transformation ${\bf T}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$ is obtained by numerically integrating the desired twist in (9). As for the desired rotation ${\bf R}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)$, it is
@@
where ${\bf R}_{\mathcal{R}_{ed}(k)}^{\mathcal{R}_{e}(k-1)}$ is the rotation matrix corresponding to the axis $\frac{{\bf r}}{\lVert{\bf r}\rVert}$ and angle $\vartheta$:
${\bf r}={\bf R}_{\mathcal{R}_{b}}^{\mathcal{R}_{e}}(k-1)\bigl({\boldsymbol{\omega}}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)T_{s}\bigr),\quad\vartheta=\left\lVert{\bf r}\right\rVert.$ (14)
@@
${\bf x}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\begin{bmatrix}{\bf p}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top},&{\bf q}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)^{\top}\end{bmatrix}^{\top}$. (16)
The eye-hand coordination mapping and scaling can be turned into an algorithmic procedure as described in Algorithm 1.
@@
    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k):=\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{e}}^{\mathcal{R}_{e}}(k)\bigr)\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{H}_{b}}^{\mathcal{H}_{e}}(k)\bigr){\bf v}_{\mathcal{H}_{e}}^{\mathcal{H}_{b}}(k)$
@@
    ${\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{b}}(k)=\bigl({\bf I}_{2}\otimes{\bf R}_{\mathcal{R}_{e}}^{\mathcal{R}_{b}}(k)\bigr)\bigl({\boldsymbol{\Lambda}}{\bf v}_{\mathcal{R}_{ed}}^{\mathcal{R}_{e}}(k)\bigr)$